a certain quantity of gas occupies 10 liter at 25 degree Celsius and a pressure of 1 atm . at what pressure will the volume of gas be 1 liter
1
Expert's answer
2020-06-11T10:28:53-0400

Due to ideal gas law:

pV/T = const

p1V1/T1 = p2V2/T2

p2 = (p1*V1*T2)/(T1*V2) = 10 atm